# Language Model Preference Evaluation with Multiple Weak Evaluators

### v3

@@ Abstract
Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge. Existing works usually leverage an LLM as the judge for comparing LLMs’ output pairwisely, yet such model-based evaluator is weak evaluator due to conflicting preference, i.e., output A is better than B, B than C, but C than A, causing contradictory evaluation results. To address this, we introduce GED (Preference Graph Ensemble and Denoise), a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensemble and denoise these graphs for better, non-contradictory evaluation results. In particular, our method consists of two primary stages: aggregating evaluations into a unified graph and applying a denoising process to eliminate cyclic inconsistencies, ensuring a directed acyclic graph (DAG) structure. We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure. Extensive experiments on ten benchmarks demonstrate GED’s superiority in three applications: model ranking, response selection, and model alignment tasks. Notably, GED combines small LLM evaluators (e.g., Llama3-8B, Mistral-7B, Qwen2-7B) to outperform strong ones (e.g., Qwen2-72B), showcasing its effectiveness in enhancing evaluation reliability and improving model performance.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a). Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024; Boyeau et al., 2024; Chatzi et al., 2024). Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b). Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
@@
Figure 1: (a) A preference graph exhibiting cyclic inconsistencies (e.g., A $\succ$ B $\succ$ C $\succ$ A), which violate transitivity. (b) Empirical results showing that even advanced LLMs (e.g., GPT-4-o) exhibit significant noise in preference judgments, leading to inconsistent evaluations. (c) Overview of our proposed framework, GED, which ensembles multiple weak evaluators and applies denoising to recover a directed acyclic graph.
@@
However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b). For example, an LLM might evaluate three responses and conclude that Response A is better than Response B, Response B is better than Response C, yet paradoxically also rank Response C as better than Response A. These cyclic patterns violate the transitivity assumption of preferences established in prior work (Ouyang et al., 2022; Song et al., 2024; Hou et al., 2024; Liu et al., 2024), thereby undermining the reliability of evaluation results. We model this conflicting preference via the preference graph. Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B. The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph. This process is illustrated in Figure 1 (a). Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency. In this paper, we define an evaluator as weak if it produces cycles into the preference graph. As shown in Figure  1 (b), we evaluated 10 Llama3-70B (AI@Meta, 2024) responses on HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) using GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B (AI@Meta, 2024) as evaluators. Even with GPT-4-o, 64% of preference graphs in HumanEval and 38% in MATH contained cycles, highlighting persistent noise and the limitations of these models as weak evaluators.
@@
To address this, we propose a novel framework, GED (Preference Graph Ensemble and Denoise), to address the inconsistencies in preference graphs generated through pairwise evaluations. Our method involves two key steps: (1) ensembling multiple weak evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph. By aggregating evaluations from multiple weak evaluators, we ”average out” the noise and biases, resulting in a more robust approximation of the true preference structure. The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks. The overall process of GED is illustrated in Figure 1 (c). We provide a theoretical analysis demonstrating the soundness of GED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability.
To validate the practical efficacy of GED, we conduct extensive experiments across model ranking, response selection, and model alignment tasks, utilizing ten widely recognized benchmark datasets, including HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), GAIA (Mialon et al., 2023), LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022). In these experiments, GED consistently outperformed baseline methods. For instance, in the response selection task, GED achieved an average performance gain of 4.51% over baseline methods across multiple benchmarks. Additionally, GED demonstrated substantial gains in scenarios where combining preference graphs from small evaluators surpassed the performance of even stronger individual evaluators. Specifically, when using using Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators, GED exceeded the performance of using the Qwen2-72B in response selection task. These results highlight GED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings.
@@ 2 Related Work
Preference evaluation of LLMs.
Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013; Boyeau et al., 2024; Chatzi et al., 2024; Shankar et al., 2024; Naresh et al., 2024), which evaluates the generated text based on intrinsic properties and coherence with the context. Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b). However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
@@
The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b, 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
3 Methodologies
In this section, we begin by defining a preference graph, which serves as the foundation for representing pairwise preferences among candidates (Section 3.1). Building on this foundation, we introduce GED structured into three key stages (Section 3.2): (1) graph ensemble, where we aggregate individual preference graphs into a unified structure, (2) graph denoising, which removes cycles and inconsistencies to ensure the preference graph is acyclic, and (3) graph-to-ranking, where we extract a reliable ranking of candidates from the denoised graph. Below, we provide detailed descriptions of each step.
3.1 Preference Graph
@@
A preference graph is a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents $n$ candidates, $A\subseteq V\times V$ is a set of directed arcs indicating pairwise preferences, and $w:A\rightarrow\mathbb{R}^{+}$ assigns weights to arcs, representing preference strength.
For distinct $u,v\in V$, an arc $(u,v)\in A$ exists if $w(u,v)>0$, where the weight $w(u,v)$ aggregates individual preferences:
$w(u,v)=\sum_{i=1}^{k}(s_{i}(u,v)-s_{i}(v,u)),$
with $s_{i}(u,v)$ representing the score from the $i$-th source. The preference graph encapsulates aggregate preferences, with arc weights reflecting cumulative strength.
3.2 GED: Preference Graph Ensemble and Denoise
Our method, GED (Preference Graph Ensemble and Denoise), begins by performing graph ensemble to aggregate a set of preference graphs. It then applies graph denoising to ensure acyclicity, followed by graph-to-ranking to derive the final node ranking. The detailed steps are as follows:
Graph ensemble.
Given $k$ weighted graphs $G_{1},\dots,G_{k}$ with shared vertex set $V$, the ensemble graph $G_{E}=(V,A_{E},w_{E})$ is constructed by setting $A_{E}=\bigcup_{i=1}^{k}A_{i}$ and defining the weight $w_{E}(u,v)=\sum_{i=1}^{k}w_{i}(u,v)$ for each arc $(u,v)\in A_{E}$.
Graph denoising.
Graph denoising involves transforming the original graph $G=(V,A,w)$ into a DAG. This transformation is achieved by identifying and removing a set of arcs known as the Feedback Arc Set (FAS) (Gabow, 1995), which is a set of arcs whose removal makes the graph acyclic. The goal is to find a minimum FAS, denoted as $R^{*}(G)$, which is a set of arcs with the smallest total weights that needs to be removed to eliminate all cycles in $G$. To find this minimum FAS, we can order the vertices of $G$ in a specific sequence $s=\{v_{1},v_{2},\ldots,v_{n}\}$. This vertex sequence induces a FAS $R(s)$, consisting of all arcs that point against the direction of the sequence, i.e., arcs $v_{j}\rightarrow v_{i}$ where $j>i$. The graph denoising problem is thus reframed as finding an optimal vertex sequence $s^{*}$ that induces the minimal FAS, such that $R(s^{*})=R^{*}(G)$. This optimal sequence $s^{*}$ ensures that the total weights of arcs eliminated to achieve a DAG is minimized.
Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012). Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al. (1993). Details can be found in Appendix M.
Graph to ranking.
@@ 3.2 GED: Preference Graph Ensemble and Denoise
$\text{desc}(v)=\left|\{u\in V:v\rightarrow u\}\right|,$
@@
$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$
@@
3.3 Applications
We apply GED to three tasks: Response Selection (selecting the best response from LLM-generated candidates), Model Ranking (ranking models based on task performance), and Model Alignment (identifying the best instruction-response pairs for training).
Response selection.
In this task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, and the goal is to identify the optimal answer $ans_{q}^{*}$. Multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences among these candidates, constructing a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$.
Each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents the preferences of evaluator $a$, where $V_{q}$ corresponds to the candidates and $A_{a}$ indicates pairwise preferences with weights $w_{a}$ reflecting preference strength. This process is detailed in Appendix E. To aggregate these graphs, GED first merges them into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, removes cycles to obtain a DAG, and derives the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$. The top-ranked answer is selected as $ans_{q}^{*}$. Repeating this for all $q\in Q$ yields the final set $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$, representing consensus from multiple evaluators and ensuring high-quality responses.
Model Ranking.
The goal of model ranking is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to questions $Q=\{q_{1},\ldots,q_{t}\}$. Evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provide pairwise preferences for model outputs, constructing preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for each question $q$.
Here, each graph $G_{a}=(V_{q},A_{a},w_{a})$ represents preferences, where $V_{q}$ corresponds to the models and $w_{a}$ reflects preference strength. The detailed procedure is outlined in Appendix E. Using GED, we aggregate these graphs into a unified graph $G_{q}$, transform it into a DAG by removing cycles, and derive the ranking $\mathcal{R}_{q}$. Repeating this for all $q\in Q$ yields a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$. Finally, a ranking ensemble is applied to compute the overall ranking $\mathcal{R}^{*}$ across questions, reflecting model performance as assessed by evaluators, as detailed in Appendix N
Model Alignment.
In the model alignment task, the objective is to identify the best response $y^{*}$ for each instruction $x$ from candidate responses $\{y_{1},\ldots,y_{n}\}$. Evaluators $\mathcal{A}$ provide pairwise preferences, forming preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where $G_{a}=(V_{x},A_{a},w_{a})$ represents preferences over $\{y_{1},\ldots,y_{n}\}$. GED aggregates these graphs into a unified graph $G_{x}$, removes cycles to produce a DAG, and derives a ranking $\mathcal{R}_{x}$. The highest-ranked response is selected as $y^{*}$ for $x$. Repeating this for all $t$ instructions yields the final training set $\{(x_{1},y_{1}^{*}),\ldots,(x_{t},y_{t}^{*})\}$, reflecting consensus among evaluators.

### v4

@@ Abstract
Despite the remarkable success of Large Language Models (LLMs), evaluating their outputs’ quality regarding preference remains a critical challenge. While existing works usually leverage a strong LLM as the judge for comparing LLMs’ response pairwisely, such a single-evaluator approach is vulnerable to cyclic preference , i.e., output A is better than B, B than C, but C is better than A, causing contradictory evaluation results. To address this, we introduce PGED (Preference Graph Ensemble and Denoising , a novel approach that leverages multiple model-based evaluators to construct preference graphs, and then ensembles and denoises these graphs for acyclic, non-contradictory evaluation results. We provide theoretical guarantees for our framework, demonstrating its efficacy in recovering the ground truth preference structure. Extensive experiments on ten benchmarks demonstrate PGED’s superiority in three applications: 1) model ranking for evaluation, 2) response selection for test-time scaling, and 3) data selection for model fine-tuning. Notably, PGED combines small LLM evaluators (e.g., Llama3-8B, Mistral-7B, Qwen2-7B) to outperform strong ones (e.g., Qwen2-72B), showcasing its effectiveness in enhancing evaluation reliability and improving model performance.
@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced various areas of artificial intelligence, particularly natural language processing and decision-making (Wu et al. 2023; Li et al. 2023a). As LLMs become increasingly capable, effective evaluation becomes critical (Liu et al. 2023; Desmond et al. 2024; Siska et al. 2024; Boyeau et al. 2024; Chatzi et al. 2024). Among evaluation methods, preference evaluation plays a critical role in both evaluating and aligning LLMs (Rafailov et al. 2024; Yuan et al. 2024; Dubois et al. 2024b). A common practice is to rely on a strong LLM (e.g., GPT-4 (Achiam et al. 2023)) as the judge to conduct pairwise comparisons (Li et al. 2023b; Chen et al. 2023b; Wang et al. 2022).
@@
Figure 1: (a) A preference graph exhibiting cyclic inconsistencies (e.g., A $\succ$ B $\succ$ C $\succ$ A), which violate transitivity. (b) Empirical results showing that even advanced LLMs (e.g., GPT-4-o) exhibit significant noise in preference judgments, leading to inconsistent evaluations. (c) Overview of our proposed framework, PGED, which ensembles multiple preference evaluators and applies denoising to recover a directed acyclic graph.
@@
However, such model-based evaluator setups often lead to cyclic preferences, where inconsistent rankings emerge, for instance, preferring A over B, B over C, yet C over A (Naresh, Tulabandhula et al. 2024; Zhang, Yin, and Wan 2024). These cyclic patterns violate the transitivity assumption of preferences established in prior work (Ouyang et al. 2022; Song et al. 2024; Hou et al. 2024; Liu et al. 2024), thereby undermining the reliability of evaluation results. We model this conflicting preference using a preference graph, where nodes represent responses and directed edges indicate pairwise preferences. Cycles in such graphs (e.g., A $\succ$ B $\succ$ C $\succ$ A) reflect evaluation noise, as shown in Figure 1(a). Ideally, a preference graph should be a directed acyclic graph (DAG) to maintain consistency. Empirically, we evaluated 10 Llama3-70B (AI@Meta 2024) responses on HumanEval (Chen et al. 2021) and MATH (Hendrycks et al. 2021), using GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al. 2023), Qwen2-72B (Yang et al. 2024), and Llama3-8B as judges. Even with GPT-4-o, 64% of preference graphs in HumanEval and 38% in MATH contained cycles, (Figure 1(b)), demonstrating the persistent noise in LLM-based preference evaluations and motivating the need for a more robust approach.
@@
To address this, we propose a novel framework, PGED (Preference Graph Ensemble and Denoising). Our method involves two key steps: (1) ensembling multiple preference evaluators to mitigate noise introduced by individual evaluators and (2) applying a denoising process to the resulting preference graph. By aggregating evaluations from multiple individual evaluators, we "average out" the noise and biases, resulting in a more robust approximation of the true preference structure. The denoising step further refines this aggregated graph by removing inconsistencies, ensuring the final preference graph is more reliable for downstream tasks. The overall process of PGED is illustrated in Figure 1 (c). We provide a theoretical analysis demonstrating the soundness of PGED, showing that by treating each individual preference graph as a random perturbation of a ground truth DAG, our ensemble and denoising framework can recover the ground truth DAG with high probability. To validate the practical efficacy of PGED, we conduct ten tasks ranging from response selection, model ranking to model alignment tasks, utilizing ten widely recognized benchmark datasets. In these experiments, PGED consistently outperformed baseline methods. Additionally, PGED demonstrated substantial gains in scenarios where combining preference graphs from small evaluators surpassed the performance of even stronger individual evaluators. Specifically, when using using Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators, PGED exceeded the performance of using the Qwen2-72B in response selection task. These results highlight PGED’s ability to mitigate preference noise, improve consistency, and enhance model performance across diverse evaluation settings. Our contributions are summarized as follows:
  • We propose PGED, a novel framework that ensembles multiple preference evaluators and denoises the resulting preference graphs to produce acyclic and reliable evaluation outputs.
  • We provide theoretical guarantees that PGED can recover the ground-truth preference DAG under a reasonable noise model, thereby offering provable consistency.
  • We conduct extensive experiments on ten benchmark datasets and three key tasks, model ranking, response selection, and data selection for fine-tuning, demonstrating that PGED significantly improves evaluation robustness. Notably, PGED using small models outperforms even strong single evaluators like Qwen2-72B.
@@ 2 Related Work
Preference evaluation.
Classical preference aggregation includes Elo, which iteratively updates item ratings from pairwise outcomes (Jiang, Ren, and Lin 2023); Bradley–Terry (BT/BTL) models, which assign latent utilities to explain win probabilities (Bradley and Terry 1952; Lu and Boutilier 2011); and MergeSort-style procedures, which sort items via pairwise comparisons with $O(n\log n)$ queries (Shyamasundar 1996). LLM-based preference evaluation increasingly relies on a strong model (e.g., GPT-4) as a zero-shot, reference-free judge of weaker models’ outputs (Shen et al. 2023; Dubois et al. 2024b; Chen et al. 2023b). PRD combines peer ranking over pairwise preferences with peer discussion between LLMs to reach consensus (Li, Patel, and Du 2023a). ChatEval forms a multi-agent referee team that debates and evaluates responses for more reliable assessments (Chan et al. 2023a). However, LLM judges can induce contradictory preferences and cycles when aggregating across outputs, yielding noisy and inconsistent preference graphs (Naresh, Tulabandhula et al. 2024; Zhang, Yin, and Wan 2024). SPA (Kadekodi, McTavish, and Ustun 2025) produces partial-order rankings that abstain on disputed pairs to balance comparability and disagreement, while ContraSolver (Zhang, Yin, and Wan 2024) builds a preference graph with self-annotations and flags edges likely responsible for contradictions. In contrast, our approach ensembles multiple evaluator-induced graphs and denoises them via a weighted feedback-arc-set objective to recover an acyclic structure suitable for robust ranking.
@@
The concept of weak supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al. 2016; Zhang, Song, and Ratner 2023; Zhang et al. 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al. 2024a; Guo and Yang 2024; Tong et al. 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al. 2023; Yang, Ma, and Liu 2024). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak and Ghodratnama 2024; Lang, Sontag, and Vijayaraghavan 2024). Recent work also shows weak LLMs can rival human feedback quality, enabling scalable and cost-efficient alignment (Tao and Li 2024; Li et al. 2025).
3 PGED: Preference Graph Ensemble and Denoising
In this section, we first introduce key definitions and assumptions underlying our approach, including the formal definition of a preference graph and the assumption of preference transitivity (Section 3.1). Building on these preliminaries, we then present our proposed framework, PGED, which consists of three main steps: (1) Graph Ensemble (Section 3.2), aggregating multiple evaluators’ preference graphs into a single unified structure; (2) Graph Denoising (Section 3.3), removing cycles to ensure an acyclic, consistent preference structure; and (3) Graph-to-Ranking (Section 3.4), converting the denoised graph into a reliable candidate ranking. The overall process of PGED is illustrated in Figure 1(c).
3.1 Preliminarily
Transitivity of Preferences.
In preference evaluation, we assume that the ground-truth preference relation is transitive. That is, for any distinct $u,v,w\in V$, if $u\succ v$ and $v\succ w$, then it must hold that $u\succ w$:
$(u\succ v)\wedge(v\succ w)\Longrightarrow(u\succ w).$ (1)
This transitivity assumption ensures that pairwise comparisons can be consistently embedded into a global ranking.
Preference Graph.
@@ 3.1 Preliminarily
A preference graph is a directed graph $G_{P}=(V,A,w)$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ represents $n$ candidates, $A\subseteq V\times V$ is a set of directed arcs indicating pairwise preferences, and $w:A\rightarrow\mathbb{R}^{+}$ assigns weights to arcs, representing preference strength. For distinct $u,v\in V$, an arc $(u,v)\in A$ exists if $w(u,v)>0$.
3.2 Graph Ensemble
Given $k$ weighted preference graphs $G_{i}=(V,A_{i},w_{i})$ on a common vertex set $V$, define
$A_{E}$ $\triangleq\Big\{(u,v)\in V\times V:\ \sum_{i=1}^{k}w_{i}(u,v)>0\Big\},$ (2)
$w_{E}(u,v)$ $\triangleq\sum_{i=1}^{k}w_{i}(u,v),\quad(u,v)\in A_{E},$
where $k$ is the total number of preference sources and $w_{i}(u,v)$ is the preference result from the $i$-th source and $w_{i}(u,v)=0$ if $(u,v)\notin A_{i}$. The ensemble graph is $G_{E}=(V,A_{E},w_{E})$.
3.3 Graph Denoising
To ensure consistency in the aggregated preference graph, we apply a denoising step that transforms a potentially cyclic weighted digraph $G=(V,A,w)$ into a directed acyclic graph (DAG). We cast this as the classical weighted feedback arc set (WFAS) problem (Gabow 1995), which seeks a minimum-total-weight set of arcs whose removal renders the graph acyclic. Formally,
$R^{*}(G)\;=\;\arg\min_{R\subseteq A}\sum_{(u,v)\in R}w(u,v)$ (3)
$\text{s.t.}\quad(V,A\setminus R)\ \text{is acyclic}.$
A convenient formulation uses vertex sequencing. Given an ordering $s=\{v_{1},\ldots,v_{n}\}$, the induced feedback arc set $R(s)$ comprises all arcs that violate the order, i.e., $(v_{j}\to v_{i})$ with $j>i$. The denoising problem then reduces to finding $s^{*}$ that minimizes the total weight of backward edges. Computing the optimal ordering is NP-hard (Karp 2010), so we adopt an efficient greedy procedure tailored to weighted graphs. Our algorithm iteratively builds a total order by removing structurally informative vertices. At each step, we first peel all sinks (zero out-degree) and place them at the end of the order, then peel all sources (zero in-degree) and place them at the beginning. When neither exists, we choose a vertex $u$ maximizing the weighted degree difference
$\delta(u)\;=\;d^{+}(u)-d^{-}(u),$ (4)
where $d^{+}(u)=\sum_{(u,v)\in A}w(u,v)$ and $d^{-}(u)=\sum_{(v,u)\in A}w(v,u)$. Repeating until all vertices are removed yields a total order; the backward arcs under this order form an approximate feedback arc set $R(s)$. We then construct and return only the denoised graph $G^{\prime}=(V,A\setminus R(s),w)$, which is a subgraph of the input and is typically sparse rather than fully connected. The full procedure is summarized in Algorithm 1. Its theoretical guarantee can be seen in Appendix M.
Algorithm 1 Preference Graph Denoising for PGED
1:Input: Weighted digraph $G=(V,A,w)$
2:Output: Denoised graph $G^{\prime}=(V,A^{\prime},w)$
3:Let $A_{0}\leftarrow A$; initialize $s_{1}\leftarrow[\ ],\ s_{2}\leftarrow[\ ]$
4:while $V\neq\emptyset$ do
5:  while $\exists$ sink $u$ in $G$ do
6:   Prepend $u$ to $s_{2}$
7:   Remove $u$ and its incident arcs from $G$
8:  end while
9:  while $\exists$ source $u$ in $G$ do
10:   Append $u$ to $s_{1}$
11:   Remove $u$ and its incident arcs from $G$
12:  end while
13:  if $V=\emptyset$ then break
14:  end if
15:  Select $u=\arg\max_{v\in V}\big(d^{+}(v)-d^{-}(v)\big)$
16:  Append $u$ to $s_{1}$
17:  Remove $u$ and its incident arcs from $G$
18:end while
19:Concatenate $s=s_{1}\|s_{2}$
20:Compute $R(s)=\{(v_{j}\to v_{i})\in A_{0}:j>i\text{ in }s\}$
21:Set $A^{\prime}=A_{0}\setminus R(s)$
22:return $G^{\prime}=(V,A^{\prime},w)$
3.4 Graph to ranking
@@
$\text{desc}(v)=\left|\{u\in V:v\rightarrow u\}\right|,$ (5)
@@
$v_{1}\succ v_{2}\succ\cdots\succ v_{n}.$ (6)
@@
4 Downstream Tasks
We apply PGED to three tasks: Response Selection (selecting the best response from LLM-generated candidates), Model Ranking (ranking models based on task performance), and Model Alignment (identifying the best instruction-response pairs for training). Details on how preference graphs are constructed for each of the three settings can be found in Appendix H.
4.1 Response Selection
For each question $q\in Q$, a model $\mathcal{M}$ generates $n$ candidate answers $C_{q}=\{ans_{1},\ldots,ans_{n}\}$. A set of evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$ provides pairwise preferences over $C_{q}$. For each evaluator $a$, we construct a weighted preference graph $G_{a}^{q}=(V_{q},A_{a},w_{a})$, where $V_{q}=\{v_{1},\ldots,v_{n}\}$ indexes candidates and $w_{a}(u,v)$ accumulates the number of wins of $u$ over $v$ (ties are ignored). We then apply PGED: (i) Graph Ensemble aggregates $\{G_{a}^{q}\}_{a\in\mathcal{A}}$ into a single graph $G_{E}^{q}$ (Section 3.2); (ii) Graph Denoising removes a minimum-weight set of feedback arcs to produce a DAG $G^{\prime}_{q}$ (Section 3.3); and (iii) Graph-to-Ranking converts $G^{\prime}_{q}$ into a total order $\mathcal{R}_{q}=\{v_{1}\succ\cdots\succ v_{n}\}$ (Section 3.4). The top-ranked candidate is selected as $ans_{q}^{*}$. Repeating this for all $q\in Q$ yields $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$.
4.2 Model Ranking
Given a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ and questions $Q=\{q_{1},\ldots,q_{t}\}$, our goal is to produce a global ranking of $M$. For each question $q\in Q$ and evaluator $a\in\mathcal{A}$, we construct a weighted preference graph $G_{a}^{q}=(V_{q},A_{a},w_{a})$, where $V_{q}=\{m_{1},\ldots,m_{n}\}$ indexes model outputs $\mathcal{M}_{i}(q)$ and $w_{a}(u,v)$ accumulates wins of $u$ over $v$ (ties are ignored). We then apply PGED per question: (i) Graph Ensemble aggregates $\{G_{a}^{q}\}_{a\in\mathcal{A}}$ into $G_{E}^{q}$ (Section 3.2); (ii) Graph Denoising yields a DAG $G^{\prime}_{q}$ (Section 3.3); and (iii) Graph-to-Ranking returns a total order $\mathcal{R}_{q}$ over $V_{q}$ (Section 3.4). Finally, we aggregate $\{\mathcal{R}_{q}:q\in Q\}$ into an overall model ranking $\mathcal{R}^{*}$ using a ranking-ensemble procedure (Appendix Q).
4.3 Model Alignment
For each instruction $x\in X$, let the candidate set be $Y_{x}=\{y_{1},\ldots,y_{n}\}$. Evaluators $\mathcal{A}$ provide pairwise preferences over $Y_{x}$. For each $a\in\mathcal{A}$, we build a weighted preference graph $G_{a}^{x}=(V_{x},A_{a},w_{a})$, where $V_{x}=\{v_{1},\ldots,v_{n}\}$ indexes responses and $w_{a}(u,v)$ aggregates wins of $u$ over $v$ (ties ignored ). Applying PGED, ensemble, denoising, and graph-to-ranking, produces a total order $\mathcal{R}_{x}$; we select the top response $y_{x}^{*}$ for instruction $x$. Repeating this for all $x\in X$ yields the aligned training set $\{(x,y_{x}^{*}):x\in X\}$.